OpenStudy (anonymous):

Solve the logarithmic equation. Round to the nearest ten-thousandth if necessary. log 8x^3 = 4

OpenStudy (anonymous):

Applying base 10 exponential to both sides will cancel the log... \[\huge 10^{\log 8x^{3}}=10^4\] leaving you with \[8x^{3}=10000\]

OpenStudy (anonymous):

It should be easy to solve from there

OpenStudy (anonymous):

thanks!